# revtex4-2Repair the float

### v1

@@ I Introduction
Characterization of unknown quantum states and quantum processes is a critical task in quantum computing, quantum communication, and quantum information processing. Pauli Transfer Matrix (PTM)-based representation provides a versatile and natural formalism for representing quantum processes due to its inherent alignment with the structure of quantum systems and applicability across a broad spectrum of quantum computing tasks. Its utility extends to various fields such as gate set and quantum process tomography and the study of noisy dynamics in quantum systems [1, 2, 3, 4, 5, 6].
Quantum channels modeling noisy processes and various errors models can be systematically represented with PTMs, as several such channels are most naturally represented in the Pauli basis [7, 8]. Additionally, PTMs offer a way to represent quantum channels, which is easier to visualize than other representations, e.g. Kraus operators, Choi or Chi matrices or the conventional superoperator representation [9, 10]. PTMs are also widely used in the analysis of error correcting codes [11, 12, 9], error mitigation [13] and Hamiltonian learning [14]. This simply arises from the fact that the Pauli basis is a frequently used basis. This can be seen e.g. in measurement-based quantum computation [15] or in quantum error correction using stabilizers. Therefore the PTM is compatible also for experimental investigation of quantum channels.
@@ II.1 Quantum channels and their representations
Quantum channels can be represented in many different forms, each of which can be converted into the others. One common possibility of representing a given channel $\mathcal{E}\in\mathcal{L}(\operatorname{Mat}(2^{n}))$ is to choose an orthogonal basis $\mathcal{B}$ of $\operatorname{Mat}(2^{n})$ as well as an explicit ordering of its basis elements $\{B_{i}\}_{i=1}^{4^{n}}$ and to collect the entries $\langle B_{i},\mathcal{E}(B_{j})\rangle_{\text{F}}$, $i,j=1,\ldots,4^{n}$ in a $4^{n}\times 4^{n}$ matrix. To be more precise, this establishes an algebra-isomorphism
$\begin{split}[\,\cdot\,]_{\mathcal{B}}:\mathcal{L}(\operatorname{Mat}(2^{n}))&\rightarrow\operatorname{Mat}(4^{n})\\ \mathcal{E}&\mapsto(\langle B_{i},\mathcal{E}(B_{j})\rangle_{\text{F}})_{1\leq i,j\leq 4^{n}}.\end{split}$ (4)
@@
We may refer to this representation also as the canonical one and denote it by $(\mathcal{E})$. A second candidate is the aforementioned Pauli basis (3) and the resulting matrix – the Pauli transfer matrix (PTM) of $\mathcal{E}$ – holds the entries PTM$(\mathcal{E})_{\mathbf{s},\mathbf{t}}=\langle\sigma^{\mathbf{s}},\mathcal{E}(\sigma^{\mathbf{t}})\rangle_{\text{F}}$ in lexicographic order in both $\mathbf{s},\mathbf{t}\in\text{Q}^{n}$.
@@
where (9) is a basis-dependent version of the more general Choi-Jamiołkowski isomorphism [19, 20] (see also [21] for further discussions) which establishes that the complete positivity of $\mathcal{E}$ is equivalent to $\text{Choi}({\mathcal{E}})$ being a positive matrix. [^3] Normalizing the Choi matrix to unit trace yields the Choi state. Note that $\operatorname{Mat}(4^{n})$ has to be equipped with a different multiplication in order to establish multiplicativity of (9), turning it into an algebra-isomorphism: the link product
@@ 5       $B\coloneqq 0$;
9                   $B\mathrel{+}=\epsilon\texttt{-L/R-PTM}[t]\otimes\texttt{L/R-PTM}(\Omega^{\vphantom{(1)}}_{\text{t}\bm{\ast}})$;
@@
7             compute $\Omega^{(1)}_{\text{t}\bm{\ast}}$ and $\Omega^{(2)}_{\text{u}\bm{\ast}}$ according to (20);
8             if $\Omega^{(1)}_{\text{t}\bm{\ast}}\neq 0$ and $\Omega^{(2)}_{\text{u}\bm{\ast}}\neq 0$ then
9                   $B\mathrel{+}=\epsilon\texttt{-M-PTM}[t][u]\otimes\texttt{M-PTM}(\Omega_{\text{t}\bm{\ast}}^{(1)},\Omega_{\text{u}\bm{\ast}}^{(2)})$;
@@
Figure 2: PTM-algorithms for single-sided and sandwich-multiplication. The left/right multiplication algorithm inputs a single matrix $2^{n}\times 2^{n}$-matrix $A$. If $n=0$, i.e. $A$ already is a scalar, it is simply returned. Otherwise, $B$ is initialized as the zero matrix of dimension $2^{n}\times 2^{n}$. Subsequently, $A$’s CMWs are calculated via matrix slicing and, for each nonzero CMW $\Omega_{\text{t}\bm{\ast}}$, the tensor product of the $t$-th element in the lookup table for elementary PTMs for left/right multiplication with the result of L/R-PTM($\Omega_{\text{t}\bm{\ast}}$) – following (22) and (23) – is added to $B$. Finally, $B$ is returned. The algorithm for the sandwich multiplication has a similar structure, but inputs two $2^{n}\times 2^{n}$-matrices for the left and right multiplicand, respectively. If $n=0$, there scalar-valued product is returned. Otherwise, for all nonzero CMWs $\Omega^{(1)}_{\text{t}\bm{\ast}}$ and $\Omega^{(2)}_{\text{u}\bm{\ast}}$, all tensor products of the $(t,u)$-th element in the lookup table for elementary PTMs for sandwich multiplication with the result of M-PTM($\Omega^{(1)}_{\text{t}\bm{\ast}},\Omega^{(2)}_{\text{u}\bm{\ast}}$) – following (III.1.3) – are added to the as zero initialized result matrix, which is subsequently returned.
@@
$A=\sum_{\mathbf{t}\raisebox{0.45206pt}{$\scriptscriptstyle\in$}\text{Q}^{n}}\omega_{\mathbf{t}}\sigma^{\mathbf{t}},\quad\omega_{\mathbf{t}}=\langle\sigma^{\mathbf{t}},A\rangle_{\text{F}}$ (14)
@@
9                   $B\mathrel{+}=\frac{1}{2}\big{(}\epsilon\texttt{-C-PTM}[t]\otimes\texttt{AC-PTM}(\Omega_{\text{t}\bm{\ast}})\big{)}$;
10                   $B\mathrel{+}=\frac{1}{2}\big{(}\epsilon\texttt{-AC-PTM}[t]\otimes\texttt{C-PTM}(\Omega_{\text{t}\bm{\ast}})\big{)}$;
@@
9                   $B\mathrel{+}=\frac{1}{2}\big{(}\epsilon\texttt{-C-PTM}[t]\otimes\texttt{C-PTM}(\Omega_{\text{t}\bm{\ast}})\big{)}$;
10                   $B\mathrel{+}=\frac{1}{2}\big{(}\epsilon\texttt{-AC-PTM}[t]\otimes\texttt{AC-PTM}(\Omega_{\text{t}\bm{\ast}})\big{)}$;
@@
Figure 3: PTM-algorithms for the commutator and anticommutator superoperators. Both algorithms input a $2^{n}\times 2^{n}$-matrix $A$ which is simply returned in the case that $n=0$. Otherwise, $B$ is initialized as the zero matrix of dimension $2^{n}\times 2^{n}$ and $A$’s CMWs are calculated via matrix slicing. For each nonzero CMW $\Omega_{\text{t}\bm{\ast}}$, two terms added to $B$ which are different for the commutator and the anticommutator case, respectively. In the C-PTM algorithm, the halved tensor-product of the $t$-th element in the lookup table for elementary PTMs for commutators with the result of AC-PTM($\Omega_{\text{t}\bm{\ast}}$) and of the $t$-th element in the lookup table for the elementary PTMs for anticommutators with the result of C-PTM($\Omega_{\text{t}\bm{\ast}}$) are added to $B$. In the AC-PTM algorithm, in contrast, the halved tensor-product of the $t$-th element in the lookup table for elementary PTMs for commutators with the result of C-PTM($\Omega_{\text{t}\bm{\ast}}$) and of the $t$-th element in the lookup table for the elementary PTMs for anticommutators with the result of AC-PTM($\Omega_{\text{t}\bm{\ast}}$) are added to $B$. Finally, in both algorithms, $B$ is returned.
III Methods
III.1 Special superoperators
III.1.1 Single-sided multiplication
Analogously, right multiplication with $A$ yields
The first tensor factors can be pre-computed for all combination of Pauli matrices, yielding the elementary PTMs of left and right multiplication with a Pauli matrix (compare Table 1). The arising algorithmic procedure is in Figure 2 (L/R-PTM).
In summary, the PTM of a multiplication with $A$ can be calculated by essentially applying the TPD to $A$ and adding up the tensor products of the contributing elementary PTMs.
@@ III.1.2 Commutator and anticommutator
The appearing commutator, anticommutators, and products of Pauli matrices can be pre-computed again, forming a set of elementary PTMs (compare Table 1).
@@ III.1.3 Sandwich multiplication
Again, the first tensor factors can be pre-computed for all $4\times 4\times 4=64$ possible combinations of Pauli matrices, yielding the elementary PTMs for the sandwich multiplication (compare Table 2). A derived algorithmic primitive is shown in Figure 2 (M-PTM).
III.2 Change of channel representation
III.2.1 Canonical representation to PTM
The translation from the canonical representation to PTM corresponds to conducting two basis changes, one along the rows and one along the columns of $\text{Can}(\mathcal{E})$ with its matrix elements $\text{Can}(\mathcal{E})_{ij}=\langle E_{i},\mathcal{E}(E_{j})\rangle_{\text{F}}$. This can be easily seen by recalling that $\text{PTM}(\mathcal{E})_{\mathbf{s}\mathbf{t}}=\langle\sigma^{\mathbf{s}},\mathcal{E}(\sigma^{\mathbf{t}})\rangle_{\text{F}}$. Thereby, we have to apply a basis change from the canonical one to the Pauli basis along the rows (realized by applying TPD to the $4^{n}$-dimensional row vectors, reshaped to $2^{n}\times 2^{n}$-matrices) in order to express $\mathcal{E}(E_{j})$ in the Pauli basis, respectively. Subsequently, we apply the inverse basis change along the columns (realized by applying iTPD to the suitably reshaped column vectors) in order to translate from applying $\mathcal{E}$ to the canonical basis elements to its application to the Pauli strings. This algorithm is readily invertible by exchanging the application of TPD and iTPD. Furthermore, since Choi and Chi matrix are related by the very same basis change, the same algorithm(s) may also be applied to translate between these two representations.
@@ 7             $M_{\text{s},\text{t}}$coloneqq (s, t)-block of $M$;
8             $B\mathrel{+}=\epsilon\texttt{M-PTM}[\text{s}][\text{t}]\otimes\texttt{Chi-PTM}(M_{\text{s},\text{t}})$;
@@ 1 $B$coloneqq 00;
Figure 4: Algorithms for translating from the canonical representation, the Choi matrix, the Chi matrix, and the (generalized) Kraus representation to PTMs, respectively. All algorithms except Kraus-PTM input a $4^{n}\times 4^{n}$-matrix; Kraus-PTM accepts a list of pairs of $2^{n}\times 2^{n}$-matrices. Can-PTM first applies the TPD along the $4^{n}$ rows of the input matrix. Subsequently, the iTPD is applied along its $4^{n}$ columns. As in both cases (i)TPD only operates on a single row/column at once, both steps are highly parallelizable. Choi-PTM also employs the TPD as a subroutine: The $4^{n}\times 4^{n}$-dimensional input matrix is interpreted as a $2n$-qubit matrix and we execute the TPD on the first $n$ tensor factors, yielding $4^{n}$ partially expanded CMWs. Subsequently, we apply the TPD to each of these CMWs and store the results in the columns of the output matrix. Chi-PTM, in turn, is defined recursively: Return the input if it is a scalar. Otherwise, for all eight pairs (s, t) of Pauli indices, select the $4^{n-}\times 4^{n-1}$-dimensional sub-matrix $M_{\text{s},\text{t}}$ whose row indices start with s and column indices start with t and add the Kronecker product of the (s, t)-th elementary PTM for sandwich multiplication with Chi-PTM$(M_{\text{s},\text{t}})$ to the zero-initialized result matrix $B$. Lastly, Kraus-PTM iterates over all pairs of (generalized) Kraus operators and adds to result of M-PTM applied to the respective pair to the zero-initialized result $B$.
@@
Performing TPD on the individual CMWs then yields the Pauli components of the respective image $\text{PTM}(\mathcal{E})_{\mathbf{s}\mathbf{t}}=\langle\sigma^{\mathbf{s}},\mathcal{E}(\sigma^{\mathbf{t}})\rangle_{\text{F}}$ with $\mathbf{s}\in\text{Q}^{n}$. These are precisely the $(\mathbf{s},\mathbf{t}$)-entries of the PTM which already concludes the algorithmic primitive.

### v2

@@ I Introduction
Characterization of unknown quantum states and quantum processes is a critical task in quantum computing, quantum communication, and quantum information processing. Pauli Transfer Matrix (PTM)-based representation provides a versatile and natural formalism for representing quantum processes due to its inherent alignment with the structure of quantum systems and applicability across a broad spectrum of quantum computing tasks. Its utility extends to various fields such as quantum process tomography (see [1] for an application involving Pauli measurements) classical simulations of quantum circuits (e.g. in [2] PTMs are used to identify quantum channels as simulable), as well as randomized benchmarking [3] and compiling [4] where PTMs serve as the standard tool to analyze the effect of Pauli twirling.
Quantum channels modeling noisy processes and various errors models can be systematically represented with PTMs, as several such channels are most naturally represented in the Pauli basis (see, e.g., [5] for a thorough application). Additionally, PTMs offer a way to represent quantum channels, which is easier to visualize than other representations, e.g. Kraus operators, Choi or Chi matrices or the conventional superoperator representation (see [6] for an excellent overview). PTMs are also widely used in the analysis of error correcting codes (in [7] PTMs are specifically used to quantify the performance of fault-tolerant, logical gates), error mitigation (see [8] for the application of Pauli conjugation to mitigate coherent noise) and Hamiltonian learning ([9] discusses learning PTMs of general quantum processes and extends this technique to Hamiltonian learning). This simply arises from the fact that the Pauli basis is a frequently used basis. This can be seen e.g. in measurement-based quantum computation [10] or in quantum error correction using stabilizers. Therefore the PTM is compatible also for experimental investigation of quantum channels.
@@ II.1 Quantum channels and their representations
Quantum channels can be represented in many different forms, each of which can be converted into the others. One common possibility of representing a given channel $\mathcal{E}\in\mathcal{L}(\operatorname{Mat}(2^{n}))$ is to choose an orthogonal basis $\mathcal{B}$ of $\operatorname{Mat}(2^{n})$ as well as an explicit ordering of its basis elements $\{B_{i}\}_{i=1}^{4^{n}}$ and to collect the entries $\langle B_{i},\mathcal{E}(B_{j})\rangle$, $i,j=1,\ldots,4^{n}$ in a $4^{n}\times 4^{n}$ matrix. To be more precise, this establishes an algebra-isomorphism
$\begin{split}[\,\cdot\,]_{\mathcal{B}}:\mathcal{L}(\operatorname{Mat}(2^{n}))&\rightarrow\operatorname{Mat}(4^{n})\\ \mathcal{E}&\mapsto(\langle B_{i},\mathcal{E}(B_{j})\rangle)_{1\leq i,j\leq 4^{n}}.\end{split}$ (4)
@@
We may refer to this representation also as the canonical one and denote it by $(\mathcal{E})$. A second candidate is the aforementioned Pauli basis (3) and the resulting matrix – the Pauli transfer matrix (PTM) of $\mathcal{E}$ – holds the entries PTM$(\mathcal{E})_{\mathbf{s},\mathbf{t}}=\langle\sigma^{\mathbf{s}},\mathcal{E}(\sigma^{\mathbf{t}})\rangle$ in lexicographic order in both $\mathbf{s},\mathbf{t}\in\text{Q}^{n}$.
@@
where (9) is a basis-dependent version of the more general Choi-Jamiołkowski isomorphism, after Choi [14] and Jamiołkowski [15], (see also [16] for further discussions) which establishes that the complete positivity of $\mathcal{E}$ is equivalent to $\text{Choi}({\mathcal{E}})$ being a positive matrix. [^3] Normalizing the Choi matrix to unit trace yields the Choi state. Note that $\operatorname{Mat}(4^{n})$ has to be equipped with a different multiplication in order to establish multiplicativity of (9), turning it into an algebra-isomorphism: the link product
@@ 5       $B\coloneqq 0$;
9                   $B\mathrel{+}=\epsilon\texttt{-L/R-PTM}[\text{t}]\otimes\texttt{L/R-PTM}(\Omega^{\vphantom{(1)}}_{\text{t}\bm{\ast}})$;
[…]
                   // for left/right multiplication
@@
[…]
                   // for sandwich multiplication
@@
[…]
@@
$A=\sum_{\mathbf{t}\raisebox{0.45206pt}{$\scriptscriptstyle\in$}\text{Q}^{n}}\omega_{\mathbf{t}}\sigma^{\mathbf{t}},\quad\omega_{\mathbf{t}}=\langle\sigma^{\mathbf{t}},A\rangle$ (14)
@@
III Methods
III.1 Special superoperators
III.1.1 Single-sided multiplication
Analogously, right multiplication with $A$ yields
The first tensor factors can be pre-computed for all combination of Pauli matrices, yielding the elementary PTMs of left and right multiplication $\epsilon$-L/R-PTM with a Pauli matrix (compare Table 1). The arising algorithmic procedure is depicted in Figure 2 (L/R-PTM).
In summary, the PTM of a multiplication with $A$ can be calculated by essentially applying the TPD to $A$ and adding up the tensor products of the contributing elementary PTMs $\epsilon$-L/R-PTM.
@@
9                   $B\mathrel{+}=\frac{1}{2}\big{(}\epsilon\texttt{-C-PTM}[\text{t}]\otimes\texttt{AC-PTM}(\Omega_{\text{t}\bm{\ast}})\big{)}$;
10                   $B\mathrel{+}=\frac{1}{2}\big{(}\epsilon\texttt{-AC-PTM}[\text{t}]\otimes\texttt{C-PTM}(\Omega_{\text{t}\bm{\ast}})\big{)}$;
[…]
                   // for (anti-)commutators
@@
9                   $B\mathrel{+}=\frac{1}{2}\big{(}\epsilon\texttt{-C-PTM}[\text{t}]\otimes\texttt{C-PTM}(\Omega_{\text{t}\bm{\ast}})\big{)}$;
10                   $B\mathrel{+}=\frac{1}{2}\big{(}\epsilon\texttt{-AC-PTM}[\text{t}]\otimes\texttt{AC-PTM}(\Omega_{\text{t}\bm{\ast}})\big{)}$;
[…]
                   // for (anti-)commutators
@@
Figure 3: PTM-algorithms for the commutator and anticommutator superoperators. Both algorithms input a $2^{n}\times 2^{n}$-matrix $A$ which is simply returned in the case that $n=0$. Otherwise, $B$ is initialized as the zero matrix of dimension $2^{n}\times 2^{n}$ and $A$’s CMWs are calculated via matrix slicing. For each nonzero CMW $\Omega_{\text{t}\bm{\ast}}$, two terms added to $B$ which are different for the commutator and the anticommutator case, respectively. In the C-PTM algorithm, the halved tensor-product of the t-th element in the lookup table for elementary PTMs for commutators $\epsilon$-C-PTM with the result of AC-PTM($\Omega_{\text{t}\bm{\ast}}$) and of the t-th element in the lookup table for the elementary PTMs for anticommutators $\epsilon$-AC-PTM with the result of C-PTM($\Omega_{\text{t}\bm{\ast}}$) are added to $B$. In the AC-PTM algorithm, in contrast, the halved tensor-product of the t-th element in the lookup table for elementary PTMs for commutators $\epsilon$-C-PTM with the result of C-PTM($\Omega_{\text{t}\bm{\ast}}$) and of the t-th element in the lookup table for the elementary PTMs for anticommutators $\epsilon$-AC-PTM with the result of AC-PTM($\Omega_{\text{t}\bm{\ast}}$) are added to $B$. Finally, in both algorithms, $B$ is returned.
@@ III.1.2 Commutator and anticommutator
The appearing commutator, anticommutators, and products of Pauli matrices can be pre-computed again, forming a set of elementary PTMs $\epsilon$-C/AC-PTM (compare Table 1).
@@ III.1.3 Sandwich multiplication
Again, the first tensor factors can be pre-computed for all $4\times 4\times 4=64$ possible combinations of Pauli matrices, yielding the elementary PTMs for the sandwich multiplication $\epsilon$-M-PTM (compare Table 2). A derived algorithmic primitive is shown in Figure 2 (M-PTM).
III.2 Change of channel representation
III.2.1 Canonical representation to PTM
@@
The translation from the canonical representation to PTM corresponds to conducting two basis changes, one along the rows and one along the columns of $\text{Can}(\mathcal{E})$ with its matrix elements $\text{Can}(\mathcal{E})_{ij}=\langle E_{i},\mathcal{E}(E_{j})\rangle$. This can be easily seen by recalling that $\text{PTM}(\mathcal{E})_{\mathbf{s}\mathbf{t}}=\langle\sigma^{\mathbf{s}},\mathcal{E}(\sigma^{\mathbf{t}})\rangle$. Thereby, we have to apply a basis change from the canonical one to the Pauli basis along the rows (realized by applying TPD to the $4^{n}$-dimensional row vectors, reshaped to $2^{n}\times 2^{n}$-matrices) in order to express $\mathcal{E}(E_{j})$ in the Pauli basis, respectively. Subsequently, we apply the inverse basis change along the columns (realized by applying iTPD to the suitably reshaped column vectors) in order to translate from applying $\mathcal{E}$ to the canonical basis elements to its application to the Pauli strings. This algorithm is readily invertible by exchanging the application of TPD and iTPD. Furthermore, since Choi and Chi matrix are related by the very same basis change, the same algorithm(s) may also be applied to translate between these two representations.
@@ 8       $B[\,:\,,j]$ $=$ iTPD($W[\,:\,,j]$);
// interpret $M$ as $2n$-qubit matrix and
// apply iTPD to first $n$ tensor factors of $M$
@@ 7             $M_{\text{s},\text{t}}$coloneqq (s, t)-block of $M$ ;
             // $M_{\text{s},\text{t}}$: all rows starting with s and
             // all columns starting with t
8             $B\mathrel{+}=\epsilon\texttt{-M-PTM}[\text{s}][\text{t}]\otimes\texttt{Chi-PTM}(M_{\text{s},\text{t}})$;
[…]
             // for sandwich multiplication
@@ 1 $B$coloneqq 00;
Figure 4: Algorithms for translating from the canonical representation, the Choi matrix, the Chi matrix, and the (generalized) Kraus representation to PTMs, respectively. All algorithms except Kraus-PTM input a $4^{n}\times 4^{n}$-matrix; Kraus-PTM accepts a list of pairs of $2^{n}\times 2^{n}$-matrices. Can-PTM first applies the TPD along the $4^{n}$ rows of the input matrix. Subsequently, the iTPD is applied along its $4^{n}$ columns. As in both cases (i)TPD only operates on a single row/column at once, both steps are highly parallelizable. Choi-PTM also employs the TPD and its inverse as subroutines: The $4^{n}\times 4^{n}$-dimensional input matrix is interpreted as a $2n$-qubit matrix and we execute the iTPD on the first $n$ tensor factors, yielding $4^{n}$ partially expanded CMWs. Subsequently, we apply the TPD to each of these CMWs and store the results in the columns of the output matrix. Chi-PTM, in turn, is defined recursively: Return the input if it is a scalar. Otherwise, for all eight pairs (s, t) of Pauli indices, select the $4^{n-1}\times 4^{n-1}$-dimensional sub-matrix $M_{\text{s},\text{t}}$ whose row indices start with s and column indices start with t and add the Kronecker product of the (s, t)-th elementary PTM for sandwich multiplication with Chi-PTM$(M_{\text{s},\text{t}})$ to the zero-initialized result matrix $B$. Lastly, Kraus-PTM iterates over all pairs of (generalized) Kraus operators and adds to result of M-PTM applied to the respective pair to the zero-initialized result $B$.
@@
Performing TPD on the individual CMWs then yields the Pauli components of the respective image $\text{PTM}(\mathcal{E})_{\mathbf{s}\mathbf{t}}=\langle\sigma^{\mathbf{s}},\mathcal{E}(\sigma^{\mathbf{t}})\rangle$ with $\mathbf{s}\in\text{Q}^{n}$. These are precisely the $(\mathbf{s},\mathbf{t}$)-entries of the PTM which already concludes the algorithmic primitive.